hello and thank you this is the problem "write an equation in a standard form the passes through (-3,-5) and (4,5) if you could show the work that would be great
and 1 more it is write an equation in slope-intercept form for (0,-3),(1,1)

Answer by oberobic(2304)

m = slope = change in y / change in x
change in y = -5-5 = -10
change in x = -3-4 = -7
m = -10/-7 = 10/7
y = 10/7*x + b
Use (4,5) to find 'b'.
5 = 10/7*4 + b
5 -40/7 = b
b = -5/7
y = 10/7*x -5/7
multiply both sides by 7 to eliminate fractions
7y = 10x -5
-10x +7y = -5
.
.
m = -3-1 / 0-1 = -4/-1 = 4
y = 4x + b
Use (0,-3) to find 'b'.
-3 = 4*0 + b
b= -3
y = 4x -3
.
Done.